a letter is selected from the word statistics what is the probability the letter is an s or t
Answer #1

Case 1: Letter s is selected.

Number of favorable events = 3

Total number of events = 10

So,

Probability of selecting letter s = 3/10 = 0.3

Case 2: Letter t is selected.

Number of favorable events = 3

Total number of events = 10

So,

Probability of selecting letter t = 3/10 = 0.3

So

Probability of selecting letter s or letter t = 0.3 + 0.3 = 0.6

So,

Answer is:

0.6